OpenStudy (clara1223):

find f'(x) of f(x)=cos(1/x)+sec(8x)

zepdrix (zepdrix):

Ooo trig! :O Fun stuff!

zepdrix (zepdrix):

Do you know derivative of sec(x)? :)

OpenStudy (clara1223):

yes, tan(x)sec(x)

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sec(x)=\sec(x)\tan(x)\]\[\large\rm \frac{d}{dx}\sec(8x)=\sec(8x)\tan(8x)\cdot\color{royalblue}{(8x)'}\]Good good good. So to finish up the derivative of the second term, you need only make sure you apply your chain rule.

OpenStudy (clara1223):

so the derivative of sec(8x) is sec(8x)tan(8x)?

zepdrix (zepdrix):

Not entirely :O That's most of it. Do you see the blue portion though? You have to multiply by the derivative of the `inner function`.

zepdrix (zepdrix):

Derivative of 8x? :o

OpenStudy (clara1223):

8

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\sec(8x)=\sec(8x)\tan(8x)\cdot\color{royalblue}{(8x)'}=8\sec(8x)\tan(8x)\] Ok good! :) That takes care of the second term.

zepdrix (zepdrix):

An extra 8 popping out due to our chain rule.

OpenStudy (xapproachesinfinity):

chain rule, whenever stuff are inside think of performing that

zepdrix (zepdrix):

Remember your cosine derivative? :)

OpenStudy (clara1223):

-sin(x)

OpenStudy (xapproachesinfinity):

i mean stuff except that not of the form x only

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\cos(x)=-\sin(x)\]Good. We'll follow the same pattern, applying chain rule again though.\[\large\rm \frac{d}{dx}\cos\left(\frac{1}{x}\right)=-\sin\left(\frac{1}{x}\right)\cdot\color{royalblue}{\left(\frac{1}{x}\right)'}\]

zepdrix (zepdrix):

Do you remember how to differentiate something when it's in the denominator like that? :)

OpenStudy (clara1223):

(1/x)' = (x^-1)' = (-x^-2) = (-1/x^2)

OpenStudy (clara1223):

correct?

OpenStudy (xapproachesinfinity):

yes good!

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\cos\left(\frac{1}{x}\right)=-\sin\left(\frac{1}{x}\right)\cdot\color{orangered}{\left(-\frac{1}{x^2}\right)}\]Yayyy good job \c:/ Maybe simplify things down a tad, bring the negatives together and such.\[\large\rm \frac{d}{dx}\cos\left(\frac{1}{x}\right)=\frac{1}{x^2}\cdot\sin\left(\frac{1}{x}\right)\]

OpenStudy (clara1223):

Does that need to be simplified any further?

zepdrix (zepdrix):

No, not unless you wanted to get a common denominator between the two terms. That seems unnecessary though.

OpenStudy (anonymous):

Just thought I would add the trick to remember the derivative to those nastier trig functions: sec(x)=1/cos(x): \[cos^{-1}(x)\] and therefore by the chain rule: \[\frac{d}{dx}cos^{-1}(x)=-1*cos^{-2}(x)*(-sin(x)=tan(x)*cos^{-1}(x)=tan(x)sec(x)\]

OpenStudy (clara1223):

so the final answer is \[\frac{ 1 }{ x ^{2} }\sin(\frac{ 1 }{ x })+8\sec(8x)\tan(8x)\] ?

OpenStudy (anonymous):

you can do the same with cosec(x), tan(x), etc

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (xapproachesinfinity):

oh no there a down side on writing 1/cos that way you will get mixed up with inverse function

OpenStudy (xapproachesinfinity):

just leave it as it is!

OpenStudy (xapproachesinfinity):

or use parenthesis (cosx)^-1

OpenStudy (anonymous):

Oh yeah sorry.... implied by the context but your right poor form for those not as comfortable im sorry :/

OpenStudy (anonymous):

I usually use arc... but I get your point

OpenStudy (xapproachesinfinity):

yeah but now everyone use arc lol we use them interchangeably

OpenStudy (xapproachesinfinity):

not* everyone

OpenStudy (xapproachesinfinity):

check this derivative carla f(x)=tan(cosx)

OpenStudy (xapproachesinfinity):

see if you mastered the rule lol

OpenStudy (xapproachesinfinity):

@clara1223